OpenStudy (anonymous):

what is the answer to x squared - 2x -8 =0?

OpenStudy (yusuf):

x = -4

OpenStudy (anonymous):

factor it. (x-4)(x+2)=0 so x can equal 4 or -2

OpenStudy (anonymous):

x^2 - 2x - 8 = 0 (x-4)(x+2) = 0 x = 4 or -2 9

OpenStudy (anonymous):

Thnak You